Please help!!:
The cost C of double-dipped chocolate pretzels varies directly with the number of pounds of pretzels purchased, P. If the cost is $54.42 when 5.0 pounds are purchased, find a linear function that relates the cost C to the number of pounds of pretzels purchased P. Then find the cost C when 6.0 pounds are purchased.

Answer by Theo(13342)

the total cost of double dipped chocolate pretzels varies directly with the number of pounds of pretzels purchased.

let C = total cost of pretzels.
let P = number of pounds of pretzels bought.

Since this ia a direct proportion, we get C = k*Q where k is the constant of proportionality.

When C = 54.42, P = 5

Our equation becomes:

54.42 = k*5

divide both sides of this equation by 5 to get:

k = 10.884

if we buy 6 pounds, we can find C since we now know what the constant of proportionality is.

we get C = 10.884 * 6 = 65.304

the constant of proportionality k in this equation just happens to be the cost per pound of chocolate.

assume the cost of each pound of pretzels is 10.884.

then 5 pounds = 10.884 * 5 = 54.42

and 6 pounds = 10.884 * 6 = 65.304